The square of an integer is 97 less than the square of a consecutive integer. Find the two integers.
Answer by checkley77(12844)

(X+1)^2-X^2=97
X^2+2X+1-X^2=97
2X=97-1
2X=96
X=48 ANS. FOR THE SMALLER INTEGER.
48=1=49 ANS. FOR THE LARGER INTEGER.
PROOF:
49^2-48^2=97
2,401-2304=97